# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 5.2 Proofs Using the GLOVER
Section 7 (Back Arcs) will look into back arcs in more depth. We will prove there that when these arcs do exist, we are not able to construct a GLOVER data structure. Instead, we are able to immediately find a degree-doubling node. That is, a back arc immediately leads us to a node whose second out-neighborhood is at least as large as its first out-neighborhood. This means that back arcs do not hurt with the development of our algorithm. Instead, they only lead to the faster path to a degree-doubling node. In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node(s).
@@
GLOVER Supports Induction In any oriented graph, without back arcs, the minimum out-degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
First, remember that we are under the assumption that we do not have back arcs. We have already shown that under this assumption, the rooted neighborhoods in GLOVER form a total order. Further, they are a discrete total order. By definition, a set is considered well-ordered if every non-empty subset of that set has a least element. A discrete total order is a total order, where each element has a distinct next element in the order. Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a smallest element according to the order. This fulfills the condition of being well-ordered. Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set. Thus, the total order defined by the distance from $v_{0}$ on the set of rooted neighborhoods $R_{1},R_{2},\ldots$ constitutes a well ordering. ∎
This becomes an important concept because now, in addition to the nodes, we also have the rooted neighborhoods as subsets of $G$. These rooted neighborhoods are totally ordered and thus can also be used as a means of mathematical induction. This gives us several options of how to proceed with our proofs, but the foundation will be the GLOVER data structure, which orders the nodes based on their distance from $v_{0}$-the minimum out-degree node.
Not all oriented graphs possess the properties of being well-ordered. There is a vast body of research on these graphs that would have clearly revealed such a fundamental property over the years if it were true. In fact, counterexamples exist to show that oriented graphs can contain specific structures including back arcs that prevent a well ordering.
However, within the specific context of the Seymour Second Neighborhood Conjecture and the GLOVER data structure we have developed here, we posit that any oriented graphs that do not fit into this well-ordering immediately give rise to a degree-doubling node. This will be discussed more in Section 6 (The Decreasing Neighborhood Sequence Property) and Section 7 (Back Arcs) and how those two interact.
@@
While the SSNC might seem like a degree vs. distance problem, GLOVER doesn’t pick a side. Instead, it uses degree to establish a foundation, via the minimum out-degree node. We are then able to use that node to partition the graph, and take advantage of the ordered structure created by distance. This ordered distance is key to GLOVER’s analytical power, but it wouldn’t exist in this form without the initial degree-based choice of the anchor. Therefore, GLOVER achieves a balance by making them interdependent rather than mutually exclusive.
Having defined the Graph Level Order data structure and its key properties, we will next turn to more complex proofs. These proofs will demonstrate its power and effectiveness in addressing the SSNC. What follows are proofs that may appear to be remarkably simple but this is a direct result of the data structure, and in particular things like the total order. This data structure was constructed to capture the conjecture’s core principles. As we have already seen, this conjecture is a decades-old problem that has withstood the test of time and presented multiple hurdles to researchers. The fact that these proofs may appear simple is a testament to the power of the GLOVER data structure itself. Additionally, this total order gives a way to visually represent these proofs. This helps to easily visualize concepts like density and back arcs, hopefully providing an aha moment that fosters a deeper understanding. The GLOVER structure is designed to reveal the inherent order in rooted neighborhood expansion from the minimum out-degree node. Disruptions to this order, such as back arcs, provide key insights into why a degree-doubling node must exist.
While the Graph Level Order (GLOVER) data structure powerfully fine-tunes our intuition and aids in the discovery process, its utility must be carefully balanced with the demands of mathematical rigor. The photographs and illustrations offer an understanding of the general behavior of the nodes, but we have to be careful with that. Our mathematical proofs must not rely on this visual intuition alone. Each step must proceed from a logical step, grounded in definitions, axioms, theorems and proven results. This requires us to sometimes omit one of the very things that makes this GLOVER data structure special, its total order and how this is so easy to be represented pictorially. This is a necessary journey as GLOVER begins a transition from a simple data structure concept, to a powerful tool capable of providing the means to prove the Seymour Second Neighborhood Conjecture.
@@ 6.1 Introduction
Having laid the groundwork for the GLOVER data structure, we are ready to begin using it. The next and most important concept that we will define that will guide our research is the Decreasing Neighborhood Sequence Property (DNSP). With that and the GLOVER data structure, and the ammunition of knowing that any oriented graph without back-arcs can be represented in a GLOVER, we can assume we are working in this environment.
@@
There is a fundamental result in graph theory stating that every Directed Acyclic Graph (DAG) must contain a sync, or a vertex with no outgoing edges. This insight directly inspired the development of the Decreasing Neighborhood Sequence Property. Here, we formulate a hypergraph where each of the rooted neighborhoods $R_{0},R_{1},\ldots,R_{k}$ can now be re-interpreted as a hyper-node. The exterior arcs between rooted neighborhoods can then be viewed as hyper-arcs. This yields a hypergraph structure. When there are no back arcs, this hypergraph becomes a Hyper Directed Acyclic Graph (Hyper-DAG). As such, the rooted neighborhoods admit a total ordering from the minimum out-degree node $v_{0}$ through a terminal node that acts as a sink in the hypergraph. This provides both a theoretical foundation and a practical framework for understanding degree growth and influence propagation in oriented graphs.
@@
The DNSP formalizes what it means to investigate nodes on a local level. It is, in essence, the node-level negation of the Second Neighborhood Conjecture: a node has DNSP if its second out-neighborhood is strictly smaller than its first. Our focus is to identify what elements of a graph cause a node to have this property. We would also like to determine how many such conditions can exist before that in itself causes the load to collapse. Through this lens, we can better understand these local obstacles to the conjecture. We would like to see how far we can extend this. The following lemma tells us the methodology behind this approach. All we need is one node to fail the DNSP conditions in order for it to become a degree doubling node.
@@
In this work, we assume that the graph is oriented and unweighted without back arcs. We also assume that all nodes in this graph have the Decreasing Neighborhood Sequence Property. What that means is that every node’s second out-neighborhood is strictly smaller than its first out-neighborhood. This is still the setup for a proof by contradiction. In the previous section, we showed that any unweighted oriented graph without back arcs could be represented in a GLOVER data structure. Instead of allowing nodes to be placed anywhere across the plane, we will take advantage of this structure here. The nodes will be partitioned into rooted neighborhoods with interior arcs to the current rooted neighborhood and exterior arcs to the next rooted neighborhood. For the remainder of this paper, when we speak of oriented graphs, the underlying assumption is that it is an oriented graph in a GLOVER data structure.
@@
By definition, a node satisfies the Decreasing Neighborhood Sequence Property if $|N^{+}(x)|>|N^{++}(x)|$. If $x$ has zero interior arcs, then the size of $N^{++}(v_{0})$ would be at least $|N^{+}(v_{0})|$, violating the DNSP condition for $v_{0}$. This implies that $x$ must have at least one interior arc to another node in $R_{1}$. A node $x\in R_{1}$, will have out-neighbors in $R_{2}$, making them second out-neighbors of $v_{0}$. By definition, $v_{0}$ has $d^{+}(v_{0})$ first out-neighbors. This means that $N^{++}(v_{0})\geq N^{+}(v_{0})$.
@@
This is true for the out-neighbors of every node in $R_{1}$. This ensures that every node in $R_{1}$ has an interior degree of at least one. Hence, every node $x\in R_{1}$ satisfies $|\text{int}(v_{0},x)|\geq 1$. Thus, for all $x\in R_{1}$, $|\text{int}(v_{0},x)|\geq 1$. ∎
@@
A fundamental concept in directed graphs is that when all nodes have out-degree 1, there must be a cycle. In the context of the GLOVER data structure, this simple case provides an intuitive understanding of our load balancing mechanism. To prevent the minimum out-degree node, $v_{0}$, from becoming a degree-doubling node, the first rooted neighborhood, $R_{1}$, must exhibit internal connections with every node in $R_{1}$ participating in a cycle. This cyclic structure allows the nodes in $R_{1}$ to share the load, reducing the number of distinct out-neighbors $v_{0}$ needs at distance 2. This principle of local interconnectedness, where nodes within a rooted neighborhood form cycles to manage the flow of connections, is central to the GLOVER data structure’s ability to prevent degree doubling at each level. In particular, the presence of cycles within $R_{1}$ allows nodes to reuse neighbors as second out-neighbors of $v_{0}$, reducing the total number of distinct second out-neighbors and preventing degree doubling.
@@
Exterior Load Balancing Lemma Suppose we have an oriented graph $G$. If $x$ in rooted neighborhood $R_{i}$ has the Decreasing Neighborhood Sequence Property, then for all $y\in N^{+}(x),\mid\text{ext}(x,y)|<d^{+}(x)$. That is, for any out-neighbor $y$ of $x$, the number of exterior out-neighbors of $x$ which are reached through $y$ must be less than the out-degree of $x$.
@@
Assume, for the sake of contradiction, that there is a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$. By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$. Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$. Then:
@@
Lemma 6.3 (Exterior Load Balancing Lemma) is very similar to Lemma 6.2 (Interior Load Balancing Lemma), with the notable exception that Lemma 6.3 refers to any node with the DNSP, whereas we only made Lemma 6.2 true for the minimum out-degree node. This generalization emphasizes the pervasive role of exterior neighborhood structure in maintaining the DNSP throughout the graph.
@@
This similarity to Daamouch’s work shows the power in transitive triangles. As mentioned, other researchers have connected transitive triangles to the Seymour Second Neighborhood Conjecture. We have already mentioned the work of Brantner et al. in first bringing them to the conjecture. This paper by Daamouch did extensive work in studying transitive triangles in the context of $m$-free graphs and anti-transitive oriented graphs. He was able to prove that the conjecture holds for certain classes of these graphs-namely all classes of 5-anti-transitive graphs. He also found that these classes had at least two Seymour vertices.
Despite the overlap, this paper will be moving in a different direction from Daamouch and others with whom we share similarity. We have defined fundamental concepts like the rooted neighborhoods, the GLOVER data structure and the Decreasing Neighborhood Sequence Property, which will all help build towards a global search solution. The similarity of the proofs we have presented so far was first because this research has developed over a number of years. Second, these proofs are all building a data structure and an algorithm brick by brick. Later proofs will depend on these results. And while the work of Brantner et al., Daamouch and others is similar, it is not quite the same, as their work did not have terms like exterior neighbor defined. As such, using only references instead of explicitly showing that these proofs could be done within the GLOVER data structure would mean potentially missing a case that we could have explained.

### v12

@@ 5.2 Proofs Using Graph Level Order
Section 7 (Back Arcs) will look into back arcs in more depth. We will prove there that when these arcs do exist, we are not able to construct a Graph Level Order. Instead, we are able to immediately find a degree-doubling node. That is, a back arc immediately leads us to a node whose second out-neighborhood is at least as large as its first out-neighborhood. This means that back arcs do not hurt with the development of our algorithm. Instead, they only lead to the faster path to a degree-doubling node. In addition, there will no longer be a need to represent the oriented graph in a Graph Level Order, as we will have already found the degree-doubling node(s).
Next, we will follow up with our talk about transitive triangles. Now that we know that every oriented graph can be represented in a Graph Level Order, Theorem 5.2(Classification of Transitive Triangles in the Graph Level Order) gives every possible type of transitive triangle that can be represented in a Graph Level Order.
Theorem 5.2.
Classification of Transitive Triangles in the Graph Level Order In any oriented graph $G$ represented via a Graph Level Order rooted at a minimum out-degree node, every transitive triangle falls into exactly one of six structural types based on the position of nodes in rooted neighborhoods and the direction of arcs.
Proof.
We classify all transitive triangles by examining the configuration of the first two arcs and then determine whether a valid third arc can close the triangle transitively.
Case # Exterior # Back # Interior Triangle Possible?
1 0 0 3 Yes
2 0 1 2 No
3 0 2 1 Yes
4 1 0 2 Yes
5 1 1 1 Yes
6 1 2 0 Yes
7 2 0 1 Yes
8 2 1 0 No
Table 1: Enumeration of possible combinations of arc types forming transitive triangles.
We now describe the realizable cases:
  • Case 1 (Interior Triangle): All three nodes lie in the same rooted neighborhood $R_{i}$. Arcs: $x\to y,x\to z,y\to z\in G$. All arcs are interior.
  • Case 2 (Invalid): One back arc and two interior arcs. If both $y$ and $z$ lie in $R_{i}$ and one arc points backward, transitivity cannot hold unless $x$ is in $R_{i}$, which contradicts the rooted structure. Hence, this is not possible.
  • Case 3 (Back Arc Triangle I): Two back arcs and one interior arc. Here, both $x,y\in R_{i}$, and $z\in R_{i+1}$. All arcs are: $x\to y,z\to x,z\to y$. This forms a valid triangle.
  • Case 4 (Interior-Exterior Triangle): One interior arc and two exterior arcs. A common parent $x\in R_{i}$ maps to children $y,z\in R_{i+1}$ with $x\to y,x\to z,y\to z\in G$. Standard configuration for load balancing, defining the interior nodes.
[…]
  • Case 6 (Back Arc Triangle III): One exterior arc and two back arcs connecting three different rooted neighborhoods $R_{i}$, $R_{i+1}$ and $R_{i+2}$, e.g., $x\to y,z\to x,z\to y$, where $x\in R_{i}$ $y\in R_{i+1}$ and $z\in R_{i+2}$ forms a valid triangle.
  • Case 7 (Exterior Arcs Case 2:): Two exterior arcs and one interior arc, e.g., $x\to z,y\to z,x\to y$, where $x,y\in R_{i}$, $z\in R_{i+1}$.
  • Case 8 (Invalid): One back arc and two exterior arcs. The assumption of two exterior arcs implies $x\to y,x\to z)$, where $x\in R_{i}$ and $y,z\in R_{i+1}$. This third arc cannot be $y\to x$ or $z\to y$ since $G$ is an oriented graph. The other option that would make this a transitive triangle would be the arcs $y\to z$ or $z\to y$, neither of which are back arcs since $y$ and $z$ are in the same rooted neighborhood.
Interior Triangle Back Arc Triangle (Case I) Interior-Exterior Triangle
[…]
Back Arc Triangle (Case II) Back Arc Triangle (Case III) Exterior Triangle
[…]
Table 2: Examples of transitive triangle types observed in the Graph Level Order structure: Nodes are given identifiers based on their neighborhoods. From there, arcs are determined as interior, exterior, and back arc depending on the two node endpoints. We see six possible transitive triangle cases. Unlike exterior and interior arcs, back arcs are drawn with dotted lines.
Since all structurally valid triangles fall into one of these categories, and we have excluded all others through case exhaustion, the theorem follows. ∎
This further illustrates the power of the Graph Level Order. More importantly, we see transitive triangles arise naturally in domains such as social media, to represent friendship recommendations which is not a Euclidean, These social media platforms are often structured around principles like triadic closure—the tendency for two individuals with a mutual friend to form a connection—leading to the frequent emergence of transitive triangles.
What Theorem 5.2 demonstrates is that transitive triangles present a non-uniform challenge to the (SSNC). There is no one-size-fits-all approach to handling them:
  • Back arc triangles must be treated using one class of structural arguments, exploiting the fact that they are going to previous, larger neighborhoods.
  • Interior triangles require a second method, often leveraging cycles, load balance and degree-doubling within a single neighborhood level.
  • Exterior triangles necessitate a third strategy, which would focus on traversing through the remainder of the graph.
This nuanced categorization and the ability to reason about these cases independently and precisely would not be possible without the Graph Level Order, which partitions the graph by both distance and degree in a way that preserves local structure and directional influence. The Graph Level Order thus enables the decomposition of complex global conjectures into manageable local conditions, each of which can be analyzed with tools tailored to its structural type.
The presence or absence of these triangles can reveal critical information about the graph. Crucially, by categorizing these transitive triangles into these distinct types, we are able to design targeted strategies for analyzing and responding to each case. This dissection enables the development of algorithms that move beyond treating transitive triangles as a homogeneous phenomenon. Instead, we can apply a fine-grained, type-specific reasoning to each instance.
@@
Graph Level Order Supports Induction In any oriented graph, without back arcs, the minimum out-degree node $v_{0}$ within a Graph Level Order structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
@@
First, remember that we are under the assumption that we do not have back arcs. We have already shown that under this assumption, the rooted neighborhoods in a Graph Level Order form a total order. Further, they are a discrete total order. By definition, a set is considered well-ordered if every non-empty subset of that set has a least element. A discrete total order is a total order, where each element has a distinct next element in the order. Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a smallest element according to the order. This fulfills the condition of being well-ordered. Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set. Thus, the total order defined by the distance from $v_{0}$ on the set of rooted neighborhoods $R_{1},R_{2},\ldots$ constitutes a well ordering. ∎
This becomes an important concept because now, in addition to the nodes, we also have the rooted neighborhoods as subsets of $G$. These rooted neighborhoods are totally ordered and thus can also be used as a means of mathematical induction. This gives us several options of how to proceed with our proofs, but the foundation will be the Graph Level Order, which orders the nodes based on their distance from $v_{0}$-the minimum out-degree node.
Not all oriented graphs possess the properties of being well-ordered. In fact, we have been working under the assumption that there are no back arcs precisely because they are one such exception. These back arcs prevent a well ordering, which are what we need for mathematical induction. However, under the assumption of no back arcs, we do arrive at a well ordering. The proof of Lemma 5.4 established that every element has a uniquely defined next element. Back arcs would obscure this
However, within the specific context of the SSNC and the Graph Level Order we have developed here, we posit that any oriented graphs that do not fit into this well-ordering immediately give rise to a degree-doubling node. This will be discussed more in Section 6 (The Decreasing Neighborhood Sequence Property) and Section 7 (Back Arcs) and how those two interact.
@@
While the SSNC might seem like a degree vs. distance problem, the Graph Level Order doesn’t pick a side. Instead, it uses degree to establish a foundation, via the minimum out-degree node. We are then able to use that node to partition the graph, and take advantage of the ordered structure created by distance. This ordered distance is key to the Graph Level Order’s analytical power, but it wouldn’t exist in this form without the initial degree-based choice of the anchor. Therefore, the Graph Level Order achieves a balance by making them interdependent rather than mutually exclusive.
While the Graph Level Order data structure powerfully fine-tunes our intuition and aids in the discovery process, its utility must be carefully balanced with the demands of mathematical rigor. The photographs and illustrations offer an understanding of the general behavior of the nodes, but we have to be careful with that. Our mathematical proofs must not rely on this visual intuition alone. Each step must proceed from a logical step, grounded in definitions, axioms, theorems and proven results. This requires us to sometimes omit one of the very concepts that makes this Graph Level Order special, its total order and how this is so easy to be represented pictorially. This is a necessary journey as the Graph Level Order begins a transition from a simple data structure concept, to a powerful tool capable of providing the means to prove the SSNC.
@@ 6.1 Introduction
Having laid the groundwork for the Graph Level Order, we are ready to begin using it. The next and most important concept that we will define that will guide our research is the Decreasing Neighborhood Sequence Property (DNSP). With that and the Graph Level Order, and the ammunition of knowing that any oriented graph without back-arcs can be represented in a Graph Level Order, we can assume we are working in this environment.
@@
There is a fundamental result in graph theory stating that every Directed Acyclic Graph (DAG) must contain a sync [19], or a vertex with no outgoing edges. This insight directly inspired the development of the Decreasing Neighborhood Sequence Property. Here, we formulate a hypergraph where each of the rooted neighborhoods $R_{0},R_{1},\ldots,R_{k}$ can now be re-interpreted as a hyper-node. The exterior arcs between rooted neighborhoods can then be viewed as hyper-arcs. This yields a hypergraph structure. When there are no back arcs, this hypergraph becomes a Hyper Directed Acyclic Graph (Hyper-DAG). As such, the rooted neighborhoods admit a total ordering from the minimum out-degree node $v_{0}$ through a terminal node that acts as a sink in the hypergraph. This provides both a theoretical foundation and a practical framework for understanding degree growth and influence propagation in oriented graphs.
The proof by contradiction implied that there must be some inherent organization at the level of rooted neighborhoods. From standard graph theory, we knew that any directed acyclic graph (DAG) must contain a sync. By the data structure construction, all nodes are pointing forward toward that sync and away from the root node through exterior arcs. While each individual node satisfies the Decreasing Neighborhood Sequence Property (DNSP), our interest lies in whether this local condition produces a global effect on the hypergraph level. To avoid having a node’s degree double, there must be fewer exterior neighbors forward, and thus fewer exterior arcs and hyper arcs forward to the following neighborhoods. This scarcity is what ultimately causes the sizes of neighborhoods to decrease.
@@
The DNSP formalizes what it means to investigate nodes on a local level. It is, in essence, the node-level negation of the SSNC: a node has DNSP if its second out-neighborhood is strictly smaller than its first. Our focus is to identify what elements of a graph cause a node to have this property. We would also like to determine how many such conditions can exist before that in itself causes the load to collapse. Through this lens, we can better understand these local obstacles to the conjecture. We would like to see how far we can extend this. The following lemma tells us the methodology behind this approach. All we need is one node to fail the DNSP conditions in order for it to become a degree doubling node.
@@
In this work, we assume that the graph is oriented, unweighted, and without back arcs. We also assume that all nodes in this graph have the Decreasing Neighborhood Sequence Property. What that means is that every node’s second out-neighborhood is strictly smaller than its first out-neighborhood. This is still the setup for a proof by contradiction. In the previous section, we showed that any unweighted oriented graph without back arcs could be represented in a Graph Level Order. Instead of allowing nodes to be placed anywhere across the plane, we will take advantage of this structure here. The nodes will be partitioned into rooted neighborhoods with interior arcs to the current rooted neighborhood and exterior arcs to the next rooted neighborhood. For the remainder of this paper, when we speak of oriented graphs, the underlying assumption is that it is an oriented graph in a Graph Level Order.
@@
By definition, a node satisfies the Decreasing Neighborhood Sequence Property if
$|N^{+}(x)|>|N^{++}(x)|.$
By the definition of an interior arc,
$\text{int}(v_{0},x)=N^{+}(v_{0})\cap N^{+}(x)$
A node $x\in R_{1}$, will have exterior out-neighbors in $R_{2}$, making them second out-neighbors of $v_{0}$. If $x$ has zero interior arcs, then $x$ must have $\delta=d^{+}(v_{0})$ exterior arcs. This means that the size of $v_{0}$’s second neighborhood, $|N^{++}(v_{0})|$ would be at least $|N^{+}(v_{0})|$, violating the DNSP condition for $v_{0}$. This implies that $x$ must have at least one interior arc to another node in $R_{1}$. By definition, $v_{0}$ has $d^{+}(v_{0})$ first out-neighbors. This means that
$N^{++}(v_{0})\geq N^{+}(v_{0}).$
@@
This is true for the out-neighbors of every node in $R_{1}$. This ensures that every node in $R_{1}$ has an interior degree of at least one. Hence, every node $x\in R_{1}$ satisfies
$|\text{int}(v_{0},x)|\geq 1.$
Thus, for all $x\in R_{1}$,
$|\text{int}(v_{0},x)|\geq 1.$
@@
A fundamental concept in directed graphs is that when all nodes have out-degree 1, there must be a cycle. In the context of the Graph Level Order, this simple case provides an intuitive understanding of our load balancing mechanism. To prevent the minimum out-degree node, $v_{0}$, from becoming a degree-doubling node, the first rooted neighborhood, $R_{1}$, must exhibit internal connections with every node in $R_{1}$ participating in a cycle. This cyclic structure allows the nodes in $R_{1}$ to share the load, reducing the number of distinct out-neighbors $v_{0}$ needs at distance 2. This principle of local interconnectedness, where nodes within a rooted neighborhood form cycles to manage the flow of connections, is central to the Graph Level Order’s ability to prevent degree doubling at each level. In particular, the presence of cycles within $R_{1}$ allows nodes to reuse neighbors as second out-neighbors of $v_{0}$, reducing the total number of distinct second out-neighbors and preventing degree doubling.
@@
Exterior Load Balancing Lemma Suppose we have an oriented graph $G$. If $x$ in rooted neighborhood $R_{i}$ has the Decreasing Neighborhood Sequence Property, then for all $y\in N^{+}(x),|\text{ext}(x,y)|<d^{+}(x)$. That is, for any out-neighbor $y$ of $x$, the number of exterior out-neighbors of $x$ which are reached through $y$ must be less than the out-degree of $x$.
@@
Assume, for the sake of contradiction, that there is a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$. By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$. Thus,
$|\text{ext}(x,y)|\subseteq N^{++}(x).$
Then:
@@
Lemma 6.3 (Exterior Load Balancing Lemma) is very similar to Lemma 6.2 (Interior Load Balancing Lemma), with the notable exception that Lemma 6.3 refers to any node with the DNSP, whereas we only made Lemma 6.2 true for the minimum out-degree node. This generalization emphasizes the pervasive role of exterior neighborhood structure in maintaining the DNSP throughout the graph. Lemma 6.3 may seem like a stronger lemma because it applies to more nodes, but because a node’s degree is partitioned by interior and exterior neighbors, having an exterior neighbor less than the node’s overall degree is not as strong as it seems. The stronger lemma is what will come later, Lemma 6.6 (Decreasing Exteriors), which states that the size of these exteriors decrease as we move about a path in the Graph Level Order.
@@
This similarity to Daamouch’s work shows the power in transitive triangles. As mentioned, other researchers have connected transitive triangles to the SSNC. We have already mentioned the work of Brantner et al. in first bringing them to the conjecture. This paper by Daamouch did extensive work in studying transitive triangles in the context of $m$-free graphs and anti-transitive oriented graphs. He was able to prove that the conjecture holds for certain classes of these graphs-namely all classes of 5-anti-transitive graphs. He also found that these classes had at least two Seymour vertices.
Despite the overlap, this paper will be moving in a different direction from Daamouch and others with whom we share similarity. We have defined fundamental concepts like the rooted neighborhoods, the Graph Level Order and the Decreasing Neighborhood Sequence Property, which will all help build towards a global search solution. The similarity of the proofs we have presented so far was first because this research has developed over a number of years. Second, these proofs are all building a data structure and an algorithm brick by brick. Later proofs will depend on these results. And while the work of Brantner et al., Daamouch and others is similar, it is not quite the same, as their work did not have terms like exterior neighbor or the DNSP defined. As such, using only references instead of explicitly showing that these proofs could be done within the Graph Level Order would mean potentially missing a case that we could have explained.